OpenStudy (anonymous):

Find the derivative. y = π/(9x)^2

OpenStudy (anonymous):

so would I first do the power rule on the bottom and make it 18x sooo Y=pi/18x?

ganeshie8 (ganeshie8):

i think you need to change the function to x^n form first

ganeshie8 (ganeshie8):

\[\large y = \dfrac{\pi}{(9x)^2}\]

ganeshie8 (ganeshie8):

\[\large = \dfrac{\pi}{81x^2}\]

ganeshie8 (ganeshie8):

\[\large = \dfrac{\pi}{81}x^{-2}\]

ganeshie8 (ganeshie8):

now it is ready for power rule ^^

ganeshie8 (ganeshie8):

take the derivative

OpenStudy (anonymous):

or product rule? can we use that

OpenStudy (anonymous):



OpenStudy (anonymous):

is my math really off?

OpenStudy (anonymous):